OpenStudy (anonymous):

Easy equation problem Just need to make an equation and solve for this. It takes Joey 4 hours to finish refurbishing a room and it takes Brynette 6 hours. If they were to work together, how would it take to refurbish the room?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \frac{ 1 }{ 6 } =\frac{ 1 }{ x }\] I don't know if that equation would work but the 1/4 reps Joeys amount he could finih and 1/6 reps brynettes portion. x represents the number of hours it would take them to finish it together